# MARFT: Multi-Agent Reinforcement Fine-Tuning

### v2

@@ 4.2.1 Base Algorithm
7                   generate action $a^{i}\sim p_{\pi^{i}_{\theta^{i}}}(a^{i}|\hat{o}^{i},\bm{a}^{1:i-1})$
@@ 11             $\mathcal{D}\leftarrow\mathcal{D}\cup\{\left(\hat{\bm{o}}_{t},\bm{a}_{t},r_{t},\bm{o}_{t+1}\right)\}$
For the implementation of the normalization, we can directly count the token length or word length as $Z$ and divide the sum of the log probabilities of generated tokens by $Z$.
@@ 4.2.2 Token-level Adaptations
5 Preliminary Experiments
5.1 Experimental Setup
@@
Task Environment. We constructed a task environment based on the MATH (Hendrycks et al., 2021), which contains different types of math problems. The environment is initialized by randomly sampling a (problem, answer) pair from the dataset and provides no global observations to the agents. The LaMAS takes actions $(a^{1},a^{2},\dots,a^{n})$, and the environment checks the correctness of the action $a^{k}$ tagged as "with_answer" in the agent profile. The reward is a coarse-grained binary signal: 1 for correct answers and 0 otherwise. The profiles we design for these preliminary experiments are demonstrated in Appendix B.3.
Metrics. The primary metric for evaluation is the accuracy of problem-solving, computed by strictly comparing the final answer with the ground truth using mathematical conversion tools.
Baselines. We use Qwen2.5-Coder-3B-Instruct (Yang et al., 2024a; Hui et al., 2024) as the base model. The LaMAS solves math problems in a one-step style. [^4] Our framework supports multi-step rollout and optimization. However, our experimental results indicate that LaMAS, based on a 3B model, does not display the anticipated abilities in long-horizon settings, due to its limited context window length. Future experiments will include larger base models and multi-step settings. For single-agent setups, a single "actor" agent solves the problem, while in dual LaMAS, two agents ("reasoner" and "actor") collaborate. Agent profiles and prompts are detailed in the appendix. The baselines for comparison include:
  • Single-agent frameworks:
    • Untuned single agent: Single agent inference without fine-tuning.
    • Single agent action-level PPO: Fine-tuned using standard PPO (Schulman et al., 2017).
    • Single agent token-level PPO: Fine-tuned using token-level PPO, treating each token as an action.
  • Multi-agent (dual) frameworks:
    • Untuned dual LaMAS: Dual LaMAS inference without fine-tuning.
    • Dual LaMAS action-level MARFT: Fine-tuned using action-level MARFT (Algorithm 3).
    • Dual LaMAS token-level MARFT: Fine-tuned using token-level MARFT, treating each token as an action.
5.2 Analysis and Insights
Figure 3 illustrates the training dynamics of different methods across several metrics. Below, we analyze the trends by isolating variables.
(a) Correctness of dual LaMAS.
(b) Correctness of single agent.
(c) Policy loss.
(d) Action-level critic loss.
(e) Token-level critic loss.
(f) Entropy.
Figure 3: Training curves of the preliminary experiments.
Single vs. Dual. From Figures 3(a) and 3(b), it is evident that LaMAS benefits from swarm intelligence, with both baseline and tuned performances surpassing single-agent methods. Meanwhile, MARFT demonstrates better stability compared to single-agent PPO, which collapses despite initial gains as training progresses. MARFT reaches a correctness rate of approximately 50%, which is about 5% higher than the baseline. In fact, single-agent and dual LaMAS should employ different hyperparameter configurations. Supplementary experimental results in Appendix B.2 provide an alternative hyperparameter configuration, which enables single-agent PPO to enhance its policy steadily and attain a performance level similar to that of the untuned dual LaMAS.
Action-level vs. Token-level. When contrasting dual LaMAS action-level MARFT with token-level MARFT, token-level training shows early oscillations until around 8k steps, after which it surges to $\sim 0.49$ accuracy, outperforming untuned dual LaMAS by about 5%. This is consistent with the challenges of training a token-level critic, as explained in Section 3.3, i.e., the token-level critic is initially hard to train, leading to oscillating in the early stages, which also manifests in Figure 3(e).
Exception. In Figure 3(b), single-agent PPO experiences a significant collapse after 3k steps for the action-level and 6k steps for the token-level variant. Upon examining Figure 3(f), we observe substantial surges in entropy, which lead to gradient explosion and subsequent model failure. This phenomenon is a common challenge in training LLMs using RL, often arising from the agent’s exploration of bad tokens. Although difficult to eliminate, this issue can be avoided by resuming training from a stable checkpoint and mitigated by adjusting the value of $\epsilon_{\text{low}}$ and $\epsilon_{\text{high}}$ (Yu et al., 2025; Yue et al., 2025).
@@ 5.3 Future Experiments
Hyperparameter settings have a significant impact on MARFT. We include more and better preliminary experimental results in Appendix B.2 with more hyperparameter configurations, and these experiments, and we will collate the data and update the experiments in time.
Given the preliminary nature of our experiment on the only MATH (Hendrycks et al., 2021), we outline a series of future experiments to further explore the capabilities and robustness of MARFT:
  • Multi-step MARFT on mathematical benchmarks: Expand experiments to more comprehensive mathematical problem-solving tasks and benchmarks, including MATH (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), and AIME (MAA, 2024). These datasets will provide a broader evaluation of MARFT’s performance across diverse mathematical challenges.
  • Multi-step MARFT with reflective agents on coding tasks: Introduce multi-step MARFT incorporating a "reflector" agent to handle coding tasks and benchmarks such as LiveCodeBench (Jain et al., 2024) and CodeForces [^5] https://codeforces.com/. This setup will test MARFT’s ability to manage iterative problem-solving and reflection in complex coding environments.
  • Multi-step MARFT on agentic tasks: Apply MARFT to multi-step agentic tasks and benchmarks like WebShop (Yao et al., preprint) and RoCo (Mandi et al., 2023). These environments will assess MARFT’s effectiveness in scenarios requiring extended interaction and decision-making over multiple steps.
  • Hyper-parameter sensitivity analysis: Conduct extensive experiments to explore the impact of hyperparameters, particularly as the number of agents increases. This will involve testing various hyperparameter configurations to identify optimal settings that enhance stability and performance in multi-agent systems. The hyperparameter settings for the conducted preliminary experiments are listed in Appendix B.1.
@@
Through these planned experiments, we aim to validate MARFT’s versatility and address its sensitivities, ensuring its applicability across a wider range of tasks and environments.
@@ 6 Perspectives
Powerful Capabilities for Solving Complex Agentic Tasks. Agentic tasks are significantly more complex than traditional language tasks, requiring capabilities such as autonomous decision-making, collaboration, and adaptive reasoning. While a standalone LLM-based agent may lack the breadth of skills needed to address these multifaceted challenges, MARFT empowers LaMAS to overcome these limitations. By leveraging RFT, MARFT enables the efficient decomposition of complex instructions into executable sub-tasks, distributing them among specialized agents within the LaMAS framework. This allows heterogeneous LLM-based agents to contribute their unique intelligence while dynamically sharing goals, negotiating strategies, and aligning team objectives through natural language interactions. MARFT further enhances this process by optimizing both individual agent policies and collective system performance, ensuring efficient coordination and adaptability. For example, in a logistics scenario, MARFT fine-tunes agents to collaboratively solve the task of "delivering emergency supplies" by dynamically assigning sub-tasks such as route planning, inventory allocation, and risk assessment to specialized agents, thereby maximizing the system’s overall effectiveness.
@@
Scalability and Robust Generalization. MARFT offers significant scalability. Agentic multi-agent problems are highly flexible, allowing for various ways to decompose tasks and design corresponding multi-agent systems. As the agent population, environmental complexity, and task scale increase, the key challenge is to maintain efficient and stable learning performance and coordination capabilities. MARFT not only enhances the overall performance of the multi-agent system but also increases the intelligence of each agent by enabling more dynamic role-playing. Even if the system organization changes significantly, the problem can be re-decomposed to allow agents to function within a multi-agent system similar to the one they were trained in.
@@
Enhanced Privacy and Federated Learning. In a unified multi-agent system, agents do not have access to each other’s local data. Instead, they contribute their unique intelligence to the system through their actions, thereby enhancing the system’s collective intelligence. This setup aligns with many real-world scenarios where entities are willing to share their powerful agents but are highly concerned about data privacy. While this resembles federated learning, MARFT places greater emphasis on improving system performance through better collaboration among agents.
@@
Integration with Blockchain Technology. MARFT’s decentralized nature and privacy-preserving capabilities make it a natural fit for blockchain. In blockchain systems, MARFT can enable secure and efficient multi-agent collaboration without requiring agents to share sensitive data. For example, in smart contract execution, MARFT can coordinate multiple agents to validate transactions, manage decentralized autonomous organizations (DAOs), or optimize resource allocation in decentralized finance (DeFi) platforms. The ability of MARFT to dynamically adapt to changing environments and agent populations ensures robust performance in the unpredictable and adversarial settings typical of blockchain ecosystems. Additionally, MARFT’s emphasis on preserving pretrained capabilities while adapting to new tasks aligns well with the need for agents to maintain trustworthiness and reliability in blockchain-based systems. This integration opens new avenues for enhancing transparency, security, and efficiency in blockchain applications while leveraging the collective intelligence of LaMAS.
@@ 7 Open Problems
Despite the advantages and potential of MARFT, both academia and industry face numerous challenges that impede the development of more effective MARFT algorithms.
@@
Lack of Dynamic Environments for Training. A significant challenge in advancing MARFT is the absence of well-established, easy-to-implement, and scalable dynamic interactive environments designed for solving agentic tasks. There are two primary difficulties in this regard. First, creating such environments with complex agentic tasks demands substantial engineering expertise, comparable to developing fundamental environments like VirtualHome (Puig et al., 2018) or Overcooked (Carroll et al., 2019) from scratch. Second, designing reward feedback mechanisms for multi-agent systems in highly dynamic environments with complex agentic tasks is exceedingly challenging. The reward signals can be multi-dimensional, and balancing the weights among different goals is a delicate task. Although there are several "dynamic" benchmarks available, such as the Berkeley Function Calling Leaderboard (Yan et al., 2024; Patil et al., 2023), ToolSandBox (Lu et al., 2024), and GAIA (Mialon et al., 2024), converting these into dynamic environments that support MARL training remains an unresolved issue.
@@
Low Sample Efficiency and Lack of High-Quality Synthetic Data. RL, particularly on-policy RL, is known for its low sample efficiency. Algorithms like PPO and TRPO, which guarantee monotonic improvement, require frequent switching between sampling trajectories and training. This process is especially problematic and time-consuming when applied to LLMs. Potential solutions may include adopting ideas like Dyna (Sutton, 1991) to enhance sample efficiency or developing algorithms that effectively utilize stale trajectories. Another issue stemming from low sample efficiency is the scarcity of high-quality synthetic data. Effective multi-agent trajectories need to achieve high success rates while also demonstrating efficient communication and coordination. Currently, the field lacks such data for multi-agent cold starts.
@@
Lack of an Established MARFT Framework. Developing engineering solutions that integrate LLMs and MARL is highly challenging. While there exist user-friendly, efficient, and scalable frameworks such as OpenRLHF (Hu et al., 2024) for LLM fine-tuning and MALib (Zhou et al., 2023) for MARL training, a comprehensive framework specifically designed for MARFT is notably absent. The creation of such a framework would require significantly more effort than developing either of the aforementioned frameworks individually. This is because it must effectively integrate both LLMs and MARL components, each of which presents its own set of complexities. While this article introduces a MARFT framework, it is still in its early stages and offers significant room for further development and refinement.
@@
Lack of a Unified Communication Mechanism or Protocol. Recent advancements have led to the development of numerous agent communication protocols, such as MCP (Model Context Protocol) (Anthropic, 2024), A2A (Agent-to-Agent) (Surapaneni et al., 2025), and ANP (Agent Network Protocol) (Gaowei Chang, 2025). These protocols are essential for enabling effective collaboration among agents and facilitating interactions with users. They also address the critical need for agents to contribute to shared goals while preserving their own data and knowledge, reflecting the future vision of a decentralized, privacy-preserving agent ecosystem. However, the proliferation of these protocols has resulted in a fragmented landscape, lacking a unified and efficient communication pipeline. This fragmentation may hinder the seamless integration and interoperability of diverse agents, particularly in dynamic and heterogeneous environments typical of LaMAS. The absence of a standardized communication framework not only complicates system design and scalability but also introduces inefficiencies and potential bottlenecks in real-world applications.

### v3

@@ 4.2.1 Base Algorithm
7                   generate action $a^{i}\sim p_{\pi^{i}_{\theta^{i}}}(a^{i}|\hat{o}^{i},\bm{a}^{\mathcal{N}(i)}),\;\mathcal{N}(i)=\left\{j\neq i\mid\mathcal{D}(a^{j},a^{i})=1\right\}$
@@ 11             $\mathcal{D}\leftarrow\mathcal{D}\cup\{\left(\hat{\bm{o}}_{t},\bm{a}_{t},r_{t},\bm{o}_{t+1}\right)\}$
For the implementation of action normalization, we can directly count the token length or word length as $Z$ and divide the sum of the log probabilities of generated tokens by $Z$.
@@ 4.2.2 Token-level Adaptations
5 Experiments
To verify the feasibility and effectiveness of the MARFT paradigm, we implement specific and naive instantiations [^4] MARFT itself is a new paradigm, not a specific implementation. In this section, we refer to action-level and token-level MARFT as specific naive implementations based on PPO without other dedicated techniques. that follows its general principles and conduct a series of experiments [^5] All experiments are conducted on machines equipped with NVIDIA H100 GPUs (80GB), using PyTorch 2.4 and CUDA 12.1. Each experiment utilizes only 1 GPU., including main experiments of tuning LaMAS in math problem-solving environments, and ablation experiments on optimization granularity, agent population, and value pertaining. We demonstrate the learning dynamics for further analysis. More experiments and experiment details are presented in Appendix B.
5.1 Setups
Task Environment
@@ 5.1 Setups
We constructed a task environment for MATH problem-solving. The environment is supported by specific datasets and reset by randomly sampling a (problem, answer) pair from the dataset. The LaMAS takes actions, and the environment verifier checks the correctness. The reward is a coarse-grained binary signal: 1 for correct answers and 0 otherwise.
Datasets and Benchmarks
We use MATH (Hendrycks et al., 2021) and CMATH (Wei et al., 2023) for English and Chinese math problem-solving, respectively. The environments are built using the MATH train set (7.5k entries, 7 types), the CMATH train set (600 entries), and a verifier. Evaluation is conducted in-domain and cross-domain using MATH500 (500 entries) and the CMATH (1098 entries) test set, and out-of-domain using the GSM8K (Cobbe et al., 2021) test set (1319 entries).
Base Model and LaMAS Setups
We use Qwen2.5-Coder-3B-Instruct (Yang et al., 2024a; Hui et al., 2024) as the base model. Solo represents a single-agent setup, while Duo represents a LaMAS with two agents (Reasoner $\rightarrow$ Actor) collaborating on problem-solving. Each agent is equipped with a dedicated LoRA adapter. Profile configurations for these experiments are detailed in Appendix B.3.
5.2 Learning Dynamics
Figure 4 shows the training dynamics of MARFT-tuned Duo in environments supported by MATH and CMATH. The episodic return is the correctness of solving the math problem. The average step reward is calculated by dividing the episodic return by the steps taken to solve the problem.
(a) Episodic Return (MATH).
(b) Average Step Reward (MATH).
(c) Value Loss (MATH).
(d) Episodic Return (CMATH).
(e) Average Step Reward (CMATH).
(f) Value Loss (CMATH).
Figure 4: Learning dynamics during Duo MARFT (computed over 3 random seeds).
We observe that in both the MATH and CMATH settings, the episodic return exhibits a clear improvement, up to about 8% over the baseline performance. However, during the early stages of training, the episodic return demonstrates noticeable oscillations. This instability can be explained by the behavior of the value loss curves, which display high losses during the initial training phase. Besides, we observe a constant improvement in average step reward, which means Duo LaMAS not only tries to achieve high episodic return but also learns to solve the problem more efficiently.
5.3 Evaluation Results
On top of the learning dynamics of MARFT in the environment supported by the train set, we also conduct evaluations on the test set. Table 3 presents the performance of Solo, Duo, and MARFT-tuned Duo on the MATH500, CMATH, and GSM8K test sets, including in-domain and out-of-domain assessments.
Solo vs. Duo
Comparing Solo and Duo, we observe that Duo significantly outperforms Solo on the MATH500 benchmark by approximately 7%. However, their performance is comparable on CMATH and GSM8K, likely because these datasets consist of relatively simpler problems. In such cases, a more complex solving process may hinder performance. These results suggest that a LaMAS like Duo is more effective than a single agent in solving complex mathematical problems.
Table 3: Evaluation results of Solo, Duo and MARFT-tuned Duo. Duo-MATH and Duo-CMATH mean MARFT-tuned Duo in the environments supported by MATH train set and CMATH train set.
Benchmark Solo Duo Duo-MATH Duo-CMATH
MATH500 40.4% 47.2% 49.8% 49.4%
CMATH 81.4% 81.1% 83.0% 83.2%
GSM8K 76.8% 77.3% 78.7% 79.4%
Duo vs. MARFT-Tuned Duo
On in-domain benchmarks, MARFT enhances Duo’s performance by around 2% to 3%. Additionally, MARFT improves Duo’s effectiveness in out-of-domain settings, demonstrating its capability to foster collaborative behavior within Duo. This indicates that MARFT can further augment the performance of multi-agent systems.
Out-of-Domain Evaluation
Duo-MATH and Duo-CMATH, fine-tuned on English and Chinese datasets respectively, both generalize well across domains: Duo-CMATH scores 49.4% on MATH500, and Duo-MATH reaches 83.0% on CMATH. Both also improve performance on GSM8K, indicating that MARFT enhances not only in-domain accuracy but also cross-lingual and out-of-domain generalization.
5.4 Ablation Experiments
5.4.1 Optimization Granularity
Figure 5 (Figure 7 in Appendix) shows Duo MARFT with different optimization granularity, i.e., action-level and token-level. From both subfigures, we see that action-level MARFT is overall better than token-level MARFT. Specifically, action-level MARFT achieves a higher episodic return than token-level MARFT in the end and has a more stable and faster improvement. [^6] We early stop some experiments, since as the training goes on, LaMAS may explore invalid tokens, leading to a burst of entropy and collapse of training. Currently, this is a common and unavoidable phenomenon in RL for LLMs, though it can be alleviated by using some tricks (Yu et al., 2025; Yue et al., 2025). This kind of superiority of action-level over naive token-level can be theoretically deduced by the inconsistency of their optimality, which is caused by the different MDP (Wen et al., 2024b).
(a) Episodic Return (MATH).
(b) Episodic Return (CMATH).
Figure 5: Curves of episodic return during Duo MARFT with different optimization granularity (computed over 3 random seeds).
5.4.2 Agent Number (Solo vs. Duo)
Table 4: The performance ceilings of Solo and Duo during MARFT.
Environment Solo Duo
MATH $\sim$ 45% $\sim$ 53%
CMATH $\sim$ 85% $\sim$ 87%
We also conduct MARFT with the number of agents set to 1, which reduces to single-agent PPO. The asymptotically best performances of both Solo and Duo are shown in Table 4 (Figure 8 in the Appendix), from which we observe that Solo’s peak performance reaches approximately 45% on MATH and 85% on CMATH—about 8% and 2% lower, respectively, than Duo’s during training, and the harder the problem sets are, the more obvious the differences are. This result suggests that LaMAS possesses a higher performance ceiling than a single agent, even in this relatively constrained mathematical setting, as opposed to more complex agentic scenarios.
5.4.3 Value Pretraining (Warm-up)
We investigate the impact of value pertaining, also known as warm-up, where for the first few steps, only the critic network is updated while all agents stay frozen. Figure 6 shows comparisons between MARFT with warm-up and without warm-up settings.
(a) Episodic Return (MATH).
(b) Episodic Return (CMATH).
Figure 6: Curves of episodic return during Duo MARFT with and without warm-up.
Value pretraining aims to reduce the oscillations commonly seen in the early stages of training. As shown in the figures, it mitigates the sharp initial drops caused by a randomly initialized critic. While warm-up stabilizes early learning, it also slows performance improvement. Since reinforcement learning depends on exploration and exploitation, an overly long warm-up can lead to overfitting on a narrow and limited distribution of observations. Overall, its impact is modest and largely determined by the number of warm-up steps.
@@ 5.5 Future Experiments
Hyperparameter settings have a significant impact on MARFT. We include more experimental results in Appendix B.1 with more hyperparameter configurations, and these experiments, and we will collate the data and update the experiments in time.
Given the preliminary nature of our experiment on the only MATH (Hendrycks et al., 2021) and CMATH (Wei et al., 2023), we outline a series of future experiments to further explore the capabilities and robustness of MARFT:
  • MARFT on coding tasks: Introduce MARFT to handle coding tasks and benchmarks such as LiveCodeBench (Jain et al., 2024) and CodeForces [^7] https://codeforces.com/. This setup will test MARFT’s ability to manage iterative problem-solving and reflection in complex coding environments.
  • MARFT on agentic tasks: Apply MARFT to agentic tasks and benchmarks like WebShop (Yao et al., preprint) and RoCo (Mandi et al., 2023). These environments will assess MARFT’s effectiveness in scenarios requiring extended interaction and decision-making over multiple steps and turns.
  • Hyper-parameter sensitivity analysis: Conduct extensive experiments to explore the impact of hyperparameters, particularly as the number of agents increases. This will involve testing various hyperparameter configurations to identify optimal settings that enhance stability and performance in multi-agent systems. The hyperparameter settings for the conducted experiments are listed in Appendix B.2.
@@
Through these planned experiments, we aim to validate MARFT’s versatility and generalization ability, and address its sensitivities, ensuring its applicability across a wider range of tasks and environments.
@@ 6 Perspectives
Powerful Capabilities for Solving Complex Agentic Tasks
@@
Agentic tasks are significantly more complex than traditional language tasks, requiring capabilities such as autonomous decision-making, collaboration, and adaptive reasoning. While a standalone LLM-based agent may lack the breadth of skills needed to address these multifaceted challenges, MARFT empowers LaMAS to overcome these limitations. By leveraging RFT, MARFT enables the efficient decomposition of complex instructions into executable sub-tasks, distributing them among specialized agents within the LaMAS framework. This allows heterogeneous LLM-based agents to contribute their unique intelligence while dynamically sharing goals, negotiating strategies, and aligning team objectives through natural language interactions. MARFT further enhances this process by optimizing both individual agent policies and collective system performance, ensuring efficient coordination and adaptability. For example, in a logistics scenario, MARFT fine-tunes agents to collaboratively solve the task of "delivering emergency supplies" by dynamically assigning sub-tasks such as route planning, inventory allocation, and risk assessment to specialized agents, thereby maximizing the system’s overall effectiveness.
Scalability and Robust Generalization
@@
MARFT offers significant scalability. Agentic multi-agent problems are highly flexible, allowing for various ways to decompose tasks and design corresponding multi-agent systems. As the agent population, environmental complexity, and task scale increase, the key challenge is to maintain efficient and stable learning performance and coordination capabilities. MARFT not only enhances the overall performance of the multi-agent system but also increases the intelligence of each agent by enabling more dynamic role-playing. Even if the system organization changes significantly, the problem can be re-decomposed to allow agents to function within a multi-agent system similar to the one they were trained in.
Enhanced Privacy and Federated Learning
@@
In a unified multi-agent system, agents do not have access to each other’s local data. Instead, they contribute their unique intelligence to the system through their actions, thereby enhancing the system’s collective intelligence. This setup aligns with many real-world scenarios where entities are willing to share their powerful agents but are highly concerned about data privacy. While this resembles federated learning, MARFT places greater emphasis on improving system performance through better collaboration among agents.
Integration with Blockchain Technology
@@
MARFT’s decentralized nature and privacy-preserving capabilities make it a natural fit for blockchain. In blockchain systems, MARFT can enable secure and efficient multi-agent collaboration without requiring agents to share sensitive data. For example, in smart contract execution, MARFT can coordinate multiple agents to validate transactions, manage decentralized autonomous organizations (DAOs), or optimize resource allocation in decentralized finance (DeFi) platforms. The ability of MARFT to dynamically adapt to changing environments and agent populations ensures robust performance in the unpredictable and adversarial settings typical of blockchain ecosystems. Additionally, MARFT’s emphasis on preserving pretrained capabilities while adapting to new tasks aligns well with the need for agents to maintain trustworthiness and reliability in blockchain-based systems. This integration opens new avenues for enhancing transparency, security, and efficiency in blockchain applications while leveraging the collective intelligence of LaMAS.
@@ 7 Open Problems
Despite the advantages and potential of MARFT paradigm, both academia and industry face numerous challenges that impede the development of more effective MARFT algorithms.
Lack of Dynamic Environments for Training
@@
A significant challenge in advancing MARFT is the absence of well-established, easy-to-implement, and scalable dynamic interactive environments designed for solving agentic tasks. There are two primary difficulties in this regard. First, creating such environments with complex agentic tasks demands substantial engineering expertise, comparable to developing fundamental environments like VirtualHome (Puig et al., 2018) or Overcooked (Carroll et al., 2019) from scratch. Second, designing reward feedback mechanisms for multi-agent systems in highly dynamic environments with complex agentic tasks is exceedingly challenging. The reward signals can be multi-dimensional, and balancing the weights among different goals is a delicate task. Although there are several "dynamic" benchmarks available, such as the Berkeley Function Calling Leaderboard (Yan et al., 2024; Patil et al., 2023), ToolSandBox (Lu et al., 2024), and GAIA (Mialon et al., 2024), converting these into dynamic environments that support MARL training remains an unresolved issue.
Low Sample Efficiency and Lack of High-Quality Synthetic Data
@@
RL, particularly on-policy RL, is known for its low sample efficiency. Algorithms like PPO and TRPO, which guarantee monotonic improvement, require frequent switching between sampling trajectories and training. This process is especially problematic and time-consuming when applied to LLMs. Potential solutions may include adopting ideas like Dyna (Sutton, 1991) to enhance sample efficiency or developing algorithms that effectively utilize stale trajectories. Another issue stemming from low sample efficiency is the scarcity of high-quality synthetic data. Effective multi-agent trajectories need to achieve high success rates while also demonstrating efficient communication and coordination. Currently, the field lacks such data for multi-agent cold starts.
Lack of an Established MARFT Framework
@@
Developing engineering solutions that integrate LLMs and MARL is highly challenging. While there exist user-friendly, efficient, and scalable frameworks such as OpenRLHF (Hu et al., 2024) for LLM fine-tuning and MALib (Zhou et al., 2023) for MARL training, a comprehensive framework specifically designed for MARFT is notably absent. The creation of such a framework would require significantly more effort than developing either of the aforementioned frameworks individually. This is because it must effectively integrate both LLMs and MARL components, each of which presents its own set of complexities. While this article introduces a MARFT framework, it is still in its early stages and offers significant room for further development and refinement.
Lack of a Unified Communication Mechanism or Protocol
@@
Recent advancements have led to the development of numerous agent communication protocols (Yang et al., 2025b), such as MCP (Model Context Protocol) (Anthropic, 2024), A2A (Agent-to-Agent) (Surapaneni et al., 2025), and ANP (Agent Network Protocol)  (Gaowei Chang, 2025). These protocols are essential for enabling effective collaboration among agents and facilitating interactions with users. They also address the critical need for agents to contribute to shared goals while preserving their own data and knowledge, reflecting the future vision of a decentralized, privacy-preserving agent ecosystem. However, the proliferation of these protocols has resulted in a fragmented landscape, lacking a unified and efficient communication pipeline. This fragmentation may hinder the seamless integration and interoperability of diverse agents, particularly in dynamic and heterogeneous environments typical of LaMAS. The absence of a standardized communication framework not only complicates system design and scalability but also introduces inefficiencies and potential bottlenecks in real-world applications.